Write the vertex form of the equation with the vertex at the origin, and the focus at (0,-1)
Answer by TimothyLamb(4379)

the vertex form of quadratic:
y = a(x − h)^2 + k
where (h,k) is the vertex, and is given as (0,0)
---
vertex:
(vx, vy) given as (0,0)
vx = -b/2a = 0 therefore b = 0
vy = -D/4a = 0 therefore D = 0
NOTE: the formulas for vx and vy are well known and therefore given
D is the discriminant of the quadratic
---
focus:
(fx, fy) given as (0,-1)
fx = -b/2a = 0 therefore b = 0
fy = (1 - D)/4a
NOTE: the formulas for fx and fy are well known and therefore given
D is the discriminant of the quadratic, but from the vertex work above, D = 0, so:
fy = 1/4a = -1
---
therefore:
a = (-1/4)
---
the vertex form of quadratic, again:
y = a(x − h)^2 + k
where (h,k) is the vertex, and is given as (0,0)
---
Answer:
y = (-1/4)x^2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php